# Hypergraph Change Point Detection using Adapted Cardinality-Based Gadgets: Applications in Dynamic Legal Structures

### v1

@@ 2.1 Hypergraph Conductance and Graph Reduction
Another line of work uses graph reduction techniques to minimize hypergraph conductance [16, 23]. In [16, 23], the authors demonstrated that by replacing each hyperedge with a set of CB-gadgets, a hypergraph can be transformed into a directed graph while preserving conductance. This implies that minimizing conductance in the transformed directed graph also minimizes conductance in the original hypergraph. However, the resulting directed network includes auxiliary nodes with assumed zero degree [16], which hinders further spectral analysis.
@@
Given a hypergraph $\mathcal{H}=(V,E)$, we define a set of auxiliary nodes $V_{a}$ and $V_{b}$, along with a set of directed edges $\hat{E}$, as follows:
@@
The resulting graph $G$ has the node set $\hat{V}=V\cup V_{a}\cup V_{b}$. This adaptation does not change the cut score and $w_{e}\leq 1$; thus, $f_{e}$ becomes an all-or-nothing function:
@@
where $\beta=(\epsilon+1)\nu$. The following proposition shows that $\mu_{\mathcal{H}}(S)$ equals the harmonic mean of $\phi_{\mathcal{H}}(S)$ and $\frac{1}{\beta}\psi_{\mathcal{H}}(S)$ under reasonable conditions, which implies that it contains critical information about the hypergraph cut.
@@
$\mathcal{T}:=\{T\subset\hat{V}\mid\forall u\in T\cap(V_{a}\cup V_{b}),\,\exists u\in T\cap V,\,(u,v)\in\hat{E}\lor(v,u)\in\hat{E}\}$. (6)
@@
Additionally,
@@
Consequently, if a subset $T\subset\tau$ can be identified that achieves sufficiently low conductance on $G$, then by considering $S=T\cap V$, we can obtain a vertex set that achieves a correspondingly small $\mu_{\mathcal{H}}$. This provides a solid justification for optimizing the directed graph reduced by the adapted CB gadget.
@@ 2.2 Comparison Between Our Approach and Star Expansion
Perhaps the simplest method for performing spectral analysis on directed graphs is to use Chung’s method [18], as it is well-established for preserving cut information. Therefore, we directly apply Chung’s combinatorial Laplacian (i.e., $L:=\Phi-\frac{\Phi P+P^{\top}\Phi}{2}$), where $P$ is the transition probability matrix for a random walk on the directed graph and $\Phi$ is the stationary distribution vector. It is well known that this transformation preserves cut information, particularly for the strongly connected components [25].
@@
By further normalizing our proposed Laplacian and considering a scenario where the edge weights in the expanded or reduced graph are identical, we can prove the following. If $W=I$ we have $\Phi_{v}=\frac{1}{\mathrm{vol}(G)}\mathrm{diag}(d_{1},\ldots,d_{n})$, and $\Phi_{a}=\Phi_{b}=\frac{1}{\mathrm{vol}(G)}\mathrm{diag}(\delta(1),\ldots,\delta(m))$, allowing us to write
@@ 2.3 Laplacian Anomaly Detection
Because we aim to assess whether our approach better preserves spectral information in a dynamic setting, we use LAD [12] as our change point detection method. The basic idea behind LAD is to use the top $K$ singular values of the graph Laplacian at each time point as the embedding vector $v_{t}$ for the network. After this embedding vector is normalized, the context matrix $C:=[v_{t-l},v_{t-l+1},\dots,v_{t-1}]$ is constructed by aggregating information from past network embedding vectors, where $l$ represents the window that determines the breadth of the lookback period. Using this context matrix, we can estimate the embedding vector $\hat{v_{t}}$ for time $t$ based on past behavior. We calculate the anomaly score by comparing the actual embedding vector $v_{t}$ at time $t$ with the predicted embedding vector from the context matrix. Specifically, we compute the anomaly score $Z$ as $Z=1-v_{t}^{\top}\hat{v}_{t}$ and the change point score as $\hat{Z}_{t}=\min(Z_{t}-Z_{t-1},0)$.
@@ 3.2 Dynamic Legal Hypergraphs
(a) Histogram showing the total number of court opinions across the entire dataset.
(b) In and out degree distribution using the entire dataset.
Figure 1: Basic statistics of the U.S. court opinon dataset.
@@
Fig.1(a) displays the histogram of the total number of court opinions within the entire dataset. It is evident that the number of court opinions has generally increased, though there is a noticeable decline during the war period. Additionally, a recent drop in numbers may be attributed to the dataset not fully capturing the latest court opinions. Fig.1(b) illustrates the complementary cumulative distribution for both in-degrees and out-degrees. Notably, the distribution exhibits a power-law behavior in the upper tail, particularly for the in-degree distribution. We estimate the parameters of the power-law distribution using the methods proposed by [6, 2].
@@ 4.1 Synthetic Dataset
First, we tested whether the adapted CB-gadget better preserved the spectral properties of hypergraphs compared with clique and star expansions. For change point detection, we used LAD [12], which identifies potential change points based on an anomaly score. We set a tolerance of ±2, considering a prediction correct if the true change point was within two time points. We evaluated performance using three metrics: the F1-score when predicting $3\%$ of the time points (four change points, matching the ground truth), the average F1-score for predictions using $3\%$$15\%$ of all time points, and the average timing error when predicting $3\%$ of the time points. The timing error measures how close the predicted change points are to the actual change points by averaging the absolute differences between each true change point and its nearest prediction. The lower the timing error, the better the temporal accuracy. For the clique expansion, we used the full spectra. For the star and adapted CB-gadget expansions, we limited the auxiliary nodes to 100. We set the context window for LAD to 20.
@@ 6 Acknowledgment
This research was supported by JST’s Support for Pioneering Research Initiated by the Next Generation program (JPMJFR216Q), the UTEC-UTokyo FSI Research Grant Program, and the Grant-in-Aid for Scientific Research (KAKENHI) (JP24K03043).We thank Edanz for editing a draft of this manuscript.

### v2

@@ 2.1 Hypergraph Conductance and Graph Reduction
Another approach involves using graph reduction techniques to minimize hypergraph conductance [14, 21]. In [14], the authors demonstrate that by replacing each hyperedge with a set of CB-gadgets, a hypergraph can be transformed into a directed graph while preserving conductance. This transformation implies that minimizing conductance in the resulting directed graph also minimizes conductance in the original hypergraph. However, to fully preserve hypergraph conductance, it is necessary to assume that the auxiliary nodes have zero degree (see Theorem 3.3 of [14]), which complicates further spectral analysis of the resulting directed network, as it can no longer be treated as a strongly connected transition probability matrix of a directed network.
@@
Given a hypergraph $\mathcal{H}=(V,E)$, we define a set of auxiliary nodes $V_{a}$ and $V_{b}$, along with a set of directed edges $\hat{E}$, as follows [^1] We assign two unique auxiliary nodes to each hyperedge.:
@@
The resulting graph $G$ has the node set $\hat{V}=V\cup V_{a}\cup V_{b}$. This adaptation does not change the cut score and $w_{e}\leq 1$; thus, $f_{e}$ becomes an all-or-nothing function, as shown in [21]:
@@
where $\beta=(\epsilon+1)\nu$. We define $\epsilon=\max_{e\in E}|e|$ and $\nu=\max_{v\in S}|\{e\in E\mid v\in e\}|$. The following proposition shows that $\mu_{\mathcal{H}}(S)$ equals the harmonic mean of $\phi_{\mathcal{H}}(S)$ and $\frac{1}{\beta}\psi_{\mathcal{H}}(S)$ under reasonable conditions, which implies that it contains critical information about the hypergraph cut.
@@
$\mathcal{T}:=\{T\subset\hat{V}\mid\forall u\in T\cap(V_{a}\cup V_{b}),\,\exists v\in T\cap V,\,(u,v)\in\hat{E}\lor(v,u)\in\hat{E}\}$. (6)
@@
Recall that $\epsilon=\max_{e\in E}|e|$ and $\nu=\max_{v\in S}|\{e\in E\mid v\in e\}|$. Then we have
@@
Consequently, if a subset $T\subset\mathcal{T}$ can be identified that achieves sufficiently low conductance on $G$, then by considering $S=T\cap V$, we can obtain a vertex set that achieves a correspondingly small $\mu_{\mathcal{H}}$. This provides a solid justification for optimizing the directed graph reduced by the adapted CB gadget.
@@ 2.2 Comparison Between Our Approach and Star Expansion
Perhaps the simplest method for performing spectral analysis on directed graphs is Chung’s approach [16], which is well-established for preserving cut information. Our adaptation of the CB-gadget enables us to treat each edge weight as a transition probability. We directly apply Chung’s combinatorial Laplacian (i.e., $L:=\Phi-\frac{\Phi P+P^{\top}\Phi}{2}$), where $P$ is the transition probability matrix for a random walk on the directed graph, and $\Phi$ is a diagonal matrix containing the unique stationary distribution vector. This transformation is well known for preserving cut information, especially for strongly connected components [23].
@@
By further normalizing our proposed Laplacian and considering a scenario where the edge weights in the expanded or reduced graph are identical, we can prove the following. If $W=I$, we have ϕ= 1vol(G)(d_1, …, d_n, δ(1), …, δ(m), δ(1), …, δ(m)) as the unique stationary distribution vector:
$\phi P$ $=\frac{1}{\mathrm{vol}(G)}(\delta(1),\ldots,\delta(m))$
$=\frac{1}{\mathrm{vol}(G)}\begin{pmatrix}(\delta(1),\ldots,\delta(m))\Delta^{-1}H^{\top}\,&\,(d_{1},\ldots,d_{n})D^{-1}H\,&\,(\delta(1),\ldots,\delta(m))\end{pmatrix}$
$=\frac{1}{\mathrm{vol}(G)}\begin{pmatrix}\mathbf{1}^{\top}H^{\top}\,&\,\mathbf{1}^{\top}H\,&\,(\delta(1),\ldots,\delta(m))\end{pmatrix}$
$=\frac{1}{\mathrm{vol}(G)}(d_{1},\ldots,d_{n},\delta(1),\ldots,\delta(m),\delta(1),\ldots,\delta(m))$
This allows us to write
@@ 2.3 Laplacian Anomaly Detection
Because we aim to assess whether our approach better preserves spectral information in a dynamic setting, we use LAD [10] as our change point detection method. The basic idea behind LAD is to use the top $K$ singular values (or eigenvalues in our case) of the graph Laplacian at each time point as the embedding vector $v_{t}$ for the network. After this embedding vector is normalized, the context matrix $C:=[v_{t-l},v_{t-l+1},\dots,v_{t-1}]$ is constructed by aggregating information from past network embedding vectors, where $l$ represents the window that determines the breadth of the lookback period. Using this context matrix, we can estimate the embedding vector $\hat{v_{t}}$ for time $t$ based on past behavior. We calculate the anomaly score by comparing the actual embedding vector $v_{t}$ at time $t$ with the predicted embedding vector from the context matrix. Specifically, we compute the anomaly score $Z$ as $Z=1-v_{t}^{\top}\hat{v}_{t}$ and the change point score as $\hat{Z}_{t}=\min(Z_{t}-Z_{t-1},0)$.
@@ 4.1 Synthetic Dataset
First, we tested whether the adapted CB-gadget better preserved the spectral properties of hypergraphs compared with clique and star expansions. For change point detection, we used LAD [10], which identifies potential change points based on an anomaly score. We set a tolerance of ±2, considering a prediction correct if the true change point was within two time points. We evaluated performance using three metrics: the F1-score when predicting $3\%$ of the time points (four change points, matching the ground truth), the average F1-score for predictions using $3\%$$15\%$ of all time points, and the average timing error when predicting $3\%$ of the time points. The timing error measures how close the predicted change points are to the actual change points by averaging the absolute differences between each true change point and its nearest prediction. The lower the timing error, the better the temporal accuracy. For the clique expansion, we used the full spectra. For the star and adapted CB-gadget expansions, we limited the analysis to the top 100 eigenvalues of the spectrum. We set the context window for LAD to 20.
@@ 6 Acknowledgment
This research was supported by JST FOREST Program (Grant Number JPMJFR216Q), the UTEC-UTokyo FSI Research Grant Program, and the Grant-in-Aid for Scientific Research (KAKENHI) (JP24K03043). We thank Edanz for editing a draft of this manuscript.